OpenStudy (eric_d):

Simplify (cos5theta-cos theta)/ (sin5theta-sin theta)

OpenStudy (eric_d):

@nirmalnema

OpenStudy (eric_d):

= cos5 cos - sin5 sin is this correct so far ?

OpenStudy (eric_d):

@UnkleRhaukus

OpenStudy (eric_d):

@ganeshie8

OpenStudy (eric_d):

@paki

OpenStudy (paki):

= cos5 cos - sin5 sin ==== > how you obtained this

OpenStudy (eric_d):

yes @paki

OpenStudy (nirmalnema):

@eric_d can you simplify it further

OpenStudy (anonymous):

see this formula.you will get direct answer

OpenStudy (anonymous):

cos(a-b) = cosa cosb + sina sinb

OpenStudy (anonymous):

cos(5θ) cos(θ) + sin(5θ) sin(θ) = cos(5θ - θ) = cos(4θ)

OpenStudy (eric_d):

the given answer is -tan 3 theta

OpenStudy (nirmalnema):

@dg2 there is division not addition

OpenStudy (anonymous):

cos(5θ - θ)/sin(5θ - θ) is this the question

OpenStudy (eric_d):

(cos5theta-cos theta)/ (sin5theta-sin theta)

OpenStudy (anonymous):

same ya so we get cos 4θ/sin4θ

OpenStudy (nirmalnema):

how do u get that

OpenStudy (anonymous):

take cos as common in numerator and sin as common in denominator?

OpenStudy (nirmalnema):

that a rubbish thing u cant do that

OpenStudy (paki):

@nirmalnema be cool.. .try to lecture them regarding this please... :)

OpenStudy (anonymous):

if i made a blunder i am sorry @nirmalnema

OpenStudy (eric_d):

It's alrite........ Thanks everybody... @MM is guiding me..

OpenStudy (nirmalnema):

@dg2 its ok no problem.!! but always remember this its wrong you cant take it out common like that

OpenStudy (anonymous):

multiply cos 3θ in numertor and denominator first